# High Voltage Delivery and Distribution for the NEXT-100 Time Projection Chamber

### v2

@@ Abstract
A critical element in the realization of large liquid and gas time projection chambers (TPCs) is the delivery and distribution of high voltages into and around the detector. Such experiments require of order tens of kilovolts to enable electron drift over meter-scale distances. This paper describes the design and operation of the cathode feedthrough and high voltage distribution through the field cage of the NEXT-100 experiment, an underground TPC that will search for neutrinoless double beta decay $0\nu\beta\beta$. The feedthrough has been demonstrated to hold pressures up to 20 bar and sustain voltages as high as -65 kV, and the TPC is operating stably at its design high voltages. The system has been realized within the constraints of a stringent radiopurity budget and is now being used to execute a suite of sensitive double beta decay analyses.
@@ 1 Introduction
The Neutrino Experiment with a Xenon TPC (NEXT) program is the product of an international collaboration aiming to develop a series of high pressure gas xenon Time Projection Chambers (TPCs) to search for evidence of the Majorana nature of the neutrino through observation of neutrinoless double beta decay ($0\nu\beta\beta$). The progression of NEXT detectors has increased exponentially in scale between generations and aims to culminate in ton- to multi-ton scale modules that will execute a world-leading search for $0\nu\beta\beta$. Small to mid-scale demonstrators have included NEXT-DBDM [1] NEXT-DEMO [2] and NEXT-White [3]. These experiments have proven the detector technology capable of excellent energy resolution with 1% full-width half maximum (FWHM) at 2.5 MeV [4], along with performing a measurement of the two-neutrino double beta decay background in 136Xe [5], and a first search for $0\nu\beta\beta$ [6]. The current phase of the experiment is called NEXT-100 [7] and schematically shown in Fig. 1.
@@
Figure 1: Diagram of the NEXT-100 TPC which is held in pressurized xenon (HPGXe) with the energy readout plane at the left consisting of 60 photomultiplier tubes (PMTs) held in vacuum. The cathode is held at a high negative voltage with a buffer region behind it to safely step the voltage up from ground (GRD). The drift field is a well controlled moderate electric field that moves electrons over to the electroluminescence region where there is a high electric field that creates light amplification via electroluminescence, collected via the PMTs for an energy readout, and 3584 silicon photo multipliers (SiPMs) for the tracking readout.
@@
An advantage of gaseous xenon is that the track length of a decay particle is relatively long compared with condensed phases, around 20 cm in 13.5 bar. When a double beta decay occurs within NEXT-100, it creates an initial light signal from the excitations created when the decay particles deposit their energy in the xenon, as well as ionization electrons. These electrons can be imaged in three dimensions, and so the particle path can be reconstructed. A moderate electric field ($\cal{O}$300-400 V/cm), the drift field, draws the ionized electrons to an electroluminescence region. This region has a high electric field ($\cal{O}$10 kV/cm) where the electrons gain sufficient energy to excite the xenon atoms within their path. When an excited xenon atom de-excites it then releases a photon, resulting in an electroluminescent gain process whereby each electron can generate $10^{2}-10^{3}$ secondary scintillation photons as it crosses the EL region. The XY location of these photons is imaged via a tracking plane that consists of an array of silicon photomultipliers (SiPMs). The third dimension of the event is encoded in the time of which the electrons produce light in the EL region, which can then be converted to a Z component via the known electron drift velocity. The event energy is also recorded, both in the SiPM array, and with finer precision using a high photo-cathode coverage plane of photomultiplier tubes that count the total number of photons created. The number of photons is proportional to the initial number of electrons with remarkably low fluctuations, after correcting for losses during drift and local detector response maps. These maps are continuously produced through calibration with a circulating gaseous ${}^{83\textnormal{m}}$r source [8]. A measurement of energy with a resolution better than 1% FWHM is enabled [9] via collection of secondary scintillation from the EL region. In this way, the NEXT EL TPCs represent precise calorimetric 3D-imaging detectors. Operation of both the TPC and the EL region requires the supply and distribution of a substantial DC voltage, of order 70 kV. A 3D model of the TPC system for NEXT-100 with the relevant high voltage connections indicated is shown in Fig. 2.
@@
The drift field requirements are driven by a number of factors. When the ionization track is created, the electric field must be sufficient to prevent $e^{-}$+Xe+ recombination. The drift velocity should also be sufficiently high that the readout window associated with one event is short enough to make pileup unlikely, or to induce performance degradation of the energy resolution due to electron attachment on electronegative impurities in the gas. Uniformity of the field in the fiducial volume must be maintained in order to protect the topology of the initial track, which encodes the type of event that occurred and allows separation of signal from backgrounds [10]. Drifting electrons also undergo both transverse and longitudinal diffusion, and the diffusion scale is minimized for larger drift voltages and hence shorter drift times. As such, higher drift fields lead to more faithful images of the original ionization deposits. All of these factors drive a requirement for an electric field of order 300-400 V/cm, which is homogeneous at the sub-percent level. This field is created by a “field cage”, in which conducting rings are held at graded potentials to ensure a uniform, parallel field in the drift volume. An overview of the field cage design for NEXT-100 can be found in Section 2.
@@
Figure 2: Left: Sketch of the field cage and TPC system, with the relevant high voltages indicated. See text for details and Fig. 3 for a close-up of the components contained in the orange rectangle.
@@
In a wider context, noble element TPCs operate in a variety of phases, both liquid and gas, for numerous types of physics searches today [13, 14, 15, 16, 10]. As these experiments grow in size, the drift distances become longer, and therefore, the voltages required increase. This scaling has often resulted in TPC experiments unsuccessfully reaching their design field goals [17].
The NEXT-100 high voltage system includes three high voltage feedthroughs (HVFTs), one at the cathode generating the ionization electron drift, one at end of the field cage drift region to better control the drift field, and one at the electroluminescence region for ionization charge amplification. The system design described in this paper focuses on the cathode HVFT and is noteworthy in that it has been successful in enabling the NEXT-100 experiment to reach and hold its design high voltages following a brief commissioning period in argon and xenon gases. The successful installation along with the field cage design of the NEXT-100 experiment is discussed, including their early operation during commissioning in various gaseous media.
@@ 2 Field cage design
Figure 3: Left: Labeled cross section of the field cage centered at the cathode with the buffer region on the right and drift to the left. Right: Internal view of NEXT-100 from the energy plane side looking towards the SiPMs of the tracking plane. The gold ring is the edge of the cathode which is holding a 95% transparent mesh. Note that the copper rings and resistor boards are not visible from this angle.
@@
Attached to the field cage HDPE struts are individual panels of PTFE reflector material with 0.5 cm thickness. These reflectors slide into the struts with a dovetail joint to create a light-tube, lining the active volume to enhance photon collection efficiency. These panels are also coated in the wavelength shifter Tetraphenyl Butadiene (TPB) with a thickness of a few microns. Although the sapphire windows of NEXT-100 will transmit UV light, and the photomultiplier tubes have some sensitivity the xenon emission wavelength of 175 nm, the better PMT quantum efficiency at the TPB emission peak leads to improved energy resolution with the application of a wavelength shifter. A program of R&D has quantified the reflectivity of NEXT-100 style teflon panels vs thickness [19, 20], and it was found that 0.5 cm thick panels offer excellent reflectivity while avoiding unnecessary additional material mass. Where previous NEXT detectors have used a solid light tube, these panels were produced in sections to make coating with TPB more straightforward and controllable, and installation more manageable. Installation of one of the panels is shown in Fig. 4, d. Due to their modularity, it is anticipated that a similar design can be used for larger future detectors, following validation of performance in NEXT-100.
@@
Simulations of the electric field in the NEXT-100 field cage were done using COMSOL and can be seen in Fig. 6. These simulations assume a 2D-axis symmetric approximation has been used to evaluate the radial component of the electric field at different axial positions. The field uniformity in the drift region is constant and homogenous up to a distance of 20 mm from the PTFE walls.
@@ 3.1 Feedthrough design
The initial choice for an insulating sheath was HDPE rather than UHMW as it is the most stable while sustaining high electric fields [22]. However, our R&D showed that when cooled, it became brittle, leading to frequent cracks and leaks. UHMW was then selected as a second material choice. The coefficient of thermal expansion for UHMW polyethylene was determined empirically after the numbers in literature were found to be widely scattered [23, 24] and ultimately unreliable compared to our in-house studies. This proved to be a crucial step for achieving a good seal. Each batch of UHMW was measured before machining in order to calculate the appropriate oversizing dimensions to fit inside the grounding tube. This was accomplished by measuring the width of a block of material under various temperatures to find the rate of expansion versus temperature and sizing the cryo-fit part accordingly.
@@ 3.2 Assembly procedure
The HVFT assembly requires minimal tooling, as most part matings are accomplished using interference fits. The only tool required was a 5/16" hex rod at least 50 cm long and a threaded rod at least $\sim$30 cm in length. We also produced a special jig for holding the polyethylene sheath vertically and for ease of manipulation, which we refer to as the “installation tool”. Several systems are used for temperature control of various part to accomplish cryo-fits, including a double walled dewar filled with liquid nitrogen for cooling and a commercial sous vide immersion heater for warming. A diagram illustrating installation steps is shown in Fig. 9.
Prior to assembly, all pieces were cleaned by wiping down with ethanol and allowed to dry. The insulating HDPE sheath is placed inside a plastic bag and then submerged in a water bath warmed to $65^{\circ}$ using the immersion heater. The plastic bag keeps water from entering the polyethylene inner cavity and impeding the installation, as trapped water will not compress and will inhibit the interference fit. While the water is heating, the threaded rod is affixed to the wide end of the copper pin. After the insulator has come to temperature for 30 minutes the pin can be placed into the warmed and now expanded insulating tube from the short cone end as shown in Fig. 9. The assembly is then allowed to return to room temperature.
The grounding tube is placed securely on a base with room for the polyethylene to emerge from bottom end. The installation tool is attached to the polyethylene/copper assembly and then placed in a dewar. The dewar is filled with liquid nitrogen until at least the widest section of the insulator is covered. This will enables safe and convenient handling with cryo-safety gloves. After at least 20 minutes of dwell time the polyethylene is removed and inserted into the grounding tube, with the cold end up. If the parts have been properly sized, the UHMW contracts sufficiently to slide snugly into place against an internal lip without hammering or forcing.
Once the assembly has recovered to room temperature and the polyethylene has expanded, the installation tool is removed and the hex rod is used to insert and screw brass banana plug adapters in on both sides of the copper pin. Vacuum and hydrostatic pressure testing can be performed at this point. The conical steel cap is fabricated with excess material on the mating flange. The final operation is to machine the thickness of the flange down so the conic section of the polyethylene meets the interior conic section of the cap to provide an electrical boundary and avoid creepage current forming along the polyethylene surface. Once attached to the grounding tube, this completes the grounding layer around the insulator.
@@ 3.3 Installation and cabling
The HVFT mates to the vessel via a DN180 O-ring flange which provides both a mechanical fixture and a pressure seal. This seal also provides grounding of the outer feedthrough sheath to the vessel ground. The cable outside the vessel is stripped of its PVC layer the depth into the HVFT plus a few centimeters, while the grounding braid is a few centimeters shorter, shown in Fig. 10 Right. The external cable is then inserted through the HVFT until the banana plug is seated in the copper pin. The exposed grounding braid on the cable is then spread over the conical top of the HVFT and secured with a metallic hose clamp providing a continuous grounding shield. The fully installed assembly is shown in Fig. 15 Right.
@@ 4 External system components
The cathode voltage for NEXT-100 is supplied by a commercial high voltage DC power supply (FUG-HCP 140-100000). The cable used for the HVFT is larger than the cable accepted by the chosen high voltage power supply, and as such these two cable needed to be spliced externally to the vessel. This is done within a grounded oil-filled vessel (“splice pot”) that was modeled off a filter pot design used by the LArIAT experiment [25].
@@
Figure 12: Left: Assembly procedures for the Splice Pot. Step 1 is filling the splice pot with transformer oil. Step 2 is building the G10 tube onto the KF flange followed by Step 3 where they are attached to the adapter flange. Step 4 is the copper bar assembly where the cables will make contact. Step 5 combines the G10 tubes to the copper bar assembly, and step 6 puts it into the transformer oil. Right: NEXT-100 splice pot assembled and installed at the Laboratorio Subterráneo de Canfranc.

### v3

@@ Abstract
A critical element in the realization of large liquid and gas time projection chambers (TPCs) is the delivery and distribution of high voltages into and around the detector. Such experiments require of order tens of kilovolts to enable electron drift over meter-scale distances. This paper describes the design and operation of the cathode feedthrough and high voltage distribution through the field cage of the NEXT-100 experiment, an underground TPC that will search for neutrinoless double beta decay $0\nu\beta\beta$. The feedthrough has been demonstrated to hold pressures up to 20 bar and sustain voltages as high as -65 kV. The TPC is operating stably at its design high voltages. The system has been realized within the constraints of a stringent radiopurity budget and is now being used to execute a suite of sensitive double beta decay analyses.
@@ 1 Introduction
The Neutrino Experiment with a Xenon TPC (NEXT) program is the product of an international collaboration aiming to develop a series of high pressure gas xenon Time Projection Chambers (TPCs) to search for evidence of the Majorana nature of the neutrino through observation of neutrinoless double beta decay ($0\nu\beta\beta$). The progression of NEXT detectors has increased exponentially in scale between generations and aims to culminate in ton- to multi-ton scale modules that will execute a world-leading search for $0\nu\beta\beta$. Small to mid-scale demonstrators have included NEXT-DBDM [1] NEXT-DEMO [2] and NEXT-White [3]. These experiments have proven the detector technology capable of excellent energy resolution with 1% full-width half maximum (FWHM) at 2.5 MeV [4], along with performing a measurement of the two-neutrino double beta decay background in 136Xe [5], and a first search for $0\nu\beta\beta$ [6]. The current phase of the experiment is called NEXT-100 [7] and schematically shown in Figs. 1 and 2.
@@
Figure 1: Diagram of the NEXT-100 TPC which is held in pressurized xenon (HPGXe) with the energy readout plane at the left consisting of 60 photomultiplier tubes (PMTs) held in vacuum. The cathode is held at a high negative voltage with a buffer region behind it to safely step the voltage up from ground. The drift field is a well controlled moderate electric field that moves electrons over to the electroluminescence region where there is a high electric field that creates light amplification via electroluminescence, collected via the PMTs for an energy readout, and 3584 silicon photo multipliers (SiPMs) for the tracking readout.
Noble element TPCs operate in a variety of phases, both liquid and gas, for numerous types of physics searches today [8, 9, 10, 11, 12]. As these experiments grow in size, the drift distances become longer, and therefore, the voltages required increase. This scaling has often resulted in TPC experiments unsuccessfully reaching their design field goals [13].
@@
An advantage of gaseous xenon is that the track length of a decay particle is relatively long compared with condensed phases, around 20 cm in 13.5 bar. When a double beta decay occurs within NEXT-100, it creates an initial light signal from the excitations created when the decay particles deposit their energy in the xenon, as well as ionization electrons. These electrons can be imaged in three dimensions, and so the particle path can be reconstructed. A moderate electric field ($\cal{O}$300-400 V/cm), the drift field, draws the ionized electrons to an electroluminescence region. This region has a high electric field ($\cal{O}$10 kV/cm) where the electrons gain sufficient energy to excite the xenon atoms within their path. When an excited xenon atom de-excites it then releases a photon, resulting in an electroluminescent gain process whereby each electron can generate $10^{2}-10^{3}$ secondary scintillation photons as it crosses the electroluminescence (EL) region. The XY location of these photons is imaged via a tracking plane that consists of an array of silicon photomultipliers (SiPMs). The third dimension of the event is encoded in the time of which the electrons produce light in the EL region, which can then be converted to a Z component via the known electron drift velocity. The event energy is also recorded, both in the SiPM array, and with finer precision using a high photo-cathode coverage plane of photomultiplier tubes that count the total number of photons created. The number of photons is proportional to the initial number of electrons with remarkably low fluctuations, after correcting for losses during drift and local detector response maps. These maps are continuously produced through calibration with a circulating gaseous $^{83\textnormal{m}}$Kr source [14]. A measurement of energy with a resolution better than 1% FWHM is enabled [15] via collection of secondary scintillation from the EL region. In this way, the NEXT EL TPCs represent precise calorimetric 3D-imaging detectors. Operation of both the TPC and the EL region requires the supply and distribution of a substantial DC voltage, of order 70 kV. A 3D model of the TPC system for NEXT-100 with the relevant high voltage connections indicated is shown in Fig. 2.
@@
The drift field requirements are driven by a number of factors. When the ionization track is created, the electric field must be sufficient to prevent $e^{-}$+Xe+ recombination. The drift velocity should also be sufficiently high that the readout window associated with one event is short enough to make pileup unlikely, or to induce performance degradation of the energy resolution due to electron attachment on electronegative impurities in the gas. Uniformity of the field in the fiducial volume must be maintained in order to protect the topology of the initial track, which encodes the type of event that occurred and allows separation of signal from backgrounds [12]. Drifting electrons also undergo both transverse and longitudinal diffusion, and the diffusion scale is minimized for larger drift voltages and hence shorter drift times. As such, higher drift fields lead to more faithful images of the original ionization deposits. All of these factors drive a requirement for an electric field of order 300-400 V/cm, which is homogeneous at the sub-percent level. This field is created by a “field cage”, in which conducting rings are held at graded potentials to ensure a uniform, parallel field in the drift volume. An overview of the field cage design for NEXT-100 can be found in sec. 2.
@@
Figure 2: Left: Sketch of the field cage and TPC system, with the relevant high voltages indicated. See text for details and Fig. 3 for a close-up of the components contained in the orange rectangle. Right: Cross Section of the NEXT-100 detector design showing the field cage within a pressure vessel and connected to a high voltage feedthrough which sits on top and is explained in more detail in Section 3.
@@
NEXT-100 also includes two further high voltage feedthroughs, which provide voltages for the EL region. Since the operating parameters of this sub-system are similar to the prior iteration, NEXT-White, the previous EL region feedthroughs were re-used for these parts [3]. Supply and distribution of high voltage to the cathode and field cage in NEXT-100, however, required significantly more advanced components than used in previous NEXT experiments [2, 3]. Realizing these involved a program of design, R&D, and characterization, as well as adherence to a strict radiopurity budget, described in sec. 5. This paper focuses on development and installation of the cathode HVFT and The field cage for the NEXT-100 experiment. These components have been instrumental in enabling the rapid commissioning and operation of NEXT-100, which is reached and held its design high voltages following a brief commissioning period in argon and xenon gases. Validation, commissioning and performance of the system are discussed in sec. 6.
@@ 2 Field cage design
Figure 3: Left: Labeled cross section of the field cage centered at the cathode with the buffer region on the right and drift to the left. Dimensions are provided for the cross section of the copper rings. Right: Internal view of NEXT-100 from the energy plane side looking towards the SiPMs of the tracking plane. The gold ring is the edge of the cathode which is holding a 95% transparent mesh. Note that the copper rings and resistor boards are not visible from this angle.
@@
Attached to the field cage HDPE struts are individual panels of PTFE reflector material with 0.5 cm thickness. These reflectors slide into the struts with a dovetail joint to create a light-tube, lining the active volume to enhance photon collection efficiency. These panels are also coated in the wavelength shifter Tetraphenyl Butadiene (TPB) with a thickness of a few microns. Although the sapphire windows of NEXT-100 will transmit UV light, and the photomultiplier tubes have some sensitivity the xenon emission wavelength of 175 nm, the better PMT quantum efficiency at the TPB emission peak leads to improved energy resolution with the application of a wavelength shifter. A program of R&D has quantified the reflectivity of NEXT-100 style PTFE panels vs thickness [19], and it was found that 0.5 cm thick panels offer excellent reflectivity while avoiding unnecessary additional material mass. It is notable that this is somewhat different to the optimal thickness for liquid xenon experiments [20, 21], where PTFE reflectivity is significantly enhanced due to the high refractive index of the liquid leading to increased fresnel reflectivity [22], which is a thickness-independent effect. In the gas phase, the reflectivity is primarily due to diffusion of the photons within the PTFE volume, and it therefore exhibits a more significant thickness dependence.
Where previous NEXT detectors have used a solid light tube, these panels were produced in sections to make coating with TPB more straightforward and controllable, and installation more manageable. Installation of one of the panels is shown in Fig. 4, d. Due to their modularity, it is anticipated that a similar design can be used for larger future detectors, following validation of performance in NEXT-100.
@@
Simulations of the electric field in the NEXT-100 field cage were performed using COMSOL and can be seen in Fig. 6. These simulations use a 2D-axisymmetric approximation to evaluate the radial and axial components of the electric field. A 20 mm fiducial cut from the PTFE walls is found to be sufficient to ensure a uniform drift field for double beta decay measurements.
@@ 3.1 Feedthrough design
The initial choice for an insulating sheath was HDPE rather than UHMW as it is the most stable while sustaining high electric fields [25]. However, our R&D showed that when subjected to liquid nitrogen temperatures, it became brittle, leading to frequent cracks and leaks. UHMW was then selected as a second material choice. The coefficient of thermal expansion for UHMW polyethylene was determined empirically after the numbers in literature were found to be widely scattered [26, 27] and ultimately unreliable compared to our in-house studies. This proved to be a crucial step for achieving a good seal. Each batch of UHMW was measured before machining in order to calculate the appropriate oversizing dimensions to fit inside the grounding tube. This was accomplished by measuring the width of a block of material under various temperatures to find the rate of expansion versus temperature and sizing the cryo-fit part accordingly.
@@ 3.2 Assembly procedure
The HVFT assembly requires minimal tooling, as most part matings are accomplished using interference fits. The only tool required was a 5/16" hex rod at least 50 cm long and a threaded rod at least $\sim$30 cm in length. We also produced a special “installation jig” for holding the polyethylene sheath vertically and for ease of manipulation. Several systems are used for temperature control of various part to accomplish cryo-fits, including a double walled dewar filled with liquid nitrogen for cooling and a commercial sous vide immersion heater for warming. A diagram illustrating installation steps is shown in Fig. 9.
Prior to assembly, all pieces were cleaned by wiping down with ethanol and allowed to dry. The insulating HDPE sheath is placed inside a plastic bag and then submerged in a water bath warmed to $65^{\circ}$C using the immersion heater. The plastic bag keeps water from entering the polyethylene inner cavity and impeding the installation, as trapped water will not compress and will inhibit the interference fit. While the water is heating, the threaded rod is affixed to the wide end of the copper pin. After the insulator has come to temperature for 30 minutes the pin can be placed into the warmed and now expanded insulating tube from the short cone end as shown in Fig. 9. The assembly is then allowed to return to room temperature.
The grounding tube is placed securely on a base with room for the polyethylene to emerge from bottom end. The installation jig is attached to the polyethylene/copper assembly and then placed in a dewar. The dewar is filled with liquid nitrogen until at least the widest section of the insulator is covered. This will enables safe and convenient handling with cryo-safety gloves. After at least 20 minutes of dwell time the polyethylene is removed and inserted into the grounding tube, with the cold end up. If the parts have been properly sized, the UHMW contracts sufficiently to slide snugly into place against an internal lip without hammering or forcing.
Once the assembly has recovered to room temperature and the polyethylene has expanded, the installation jig is removed and the hex rod is used to insert and screw brass banana plug adapters in on both sides of the copper pin. Vacuum and hydrostatic pressure testing can be performed at this point. The conical steel cap is fabricated with excess material on the mating flange. The final operation is to machine the thickness of the flange down so the conic section of the polyethylene meets the interior conic section of the cap to provide an electrical boundary and avoid creepage current forming along the polyethylene surface. Once attached to the grounding tube, this completes the grounding layer around the insulator.
@@ 3.3 Installation and cabling
The HVFT mates to the vessel via a DN180 O-ring flange which provides both a mechanical fixture and a pressure seal. This flange is designed to take reusable rubber rings during prototyping stages or single-use ultra-high vacuum capable HELICOFLEX© rings during operation. Either ring type is placed in a groove on the bottom of the HVFT, creating a seal as it is tightened against the flat surface seen in Fig. 11 Left. The bolts used to tighten the HVFT to the vessel provide an immediate electrical grounding of the outer feedthrough sheath to the vessel ground. The cable outside the vessel is stripped of its PVC layer the depth into the HVFT plus a few centimeters, while the grounding braid is a few centimeters shorter, shown in Fig. 10 Right. The external cable is then inserted through the HVFT until the banana plug is seated in the copper pin. The exposed grounding braid on the cable is then spread over the conical top of the HVFT and secured with a metallic hose clamp providing a continuous grounding shield. The fully installed assembly is shown in Fig. 15 Right.
@@ 4 External system components
The cathode voltage for NEXT-100 is supplied by a commercial high voltage DC power supply (HCP 140-100000 from Fug). The cable used for the HVFT has dimensions different from the output cable of the chosen high voltage power supply, 19.3 mm vs  9.4 mm outer diameters respectively, and as such these two cables needed to be spliced externally to the vessel. This is done within a grounded oil-filled vessel (“splice pot”) that was modeled off a filter pot design used by the LArIAT experiment [28].
@@
Figure 12: Assembly procedures for the Splice Pot with a picture of the NEXT-100 splice pot assembled and installed at the Laboratorio Subterráneo de Canfranc. Step 1 is filling the splice pot with transformer oil. Step 2 is building the G10 tube onto the KF flange followed by Step 3 where they are attached to the adapter flange. Step 4 is the copper bar assembly where the cables will make contact. Step 5 combines the G10 tubes to the copper bar assembly, and step 6 puts it into the transformer oil.